### v1

# Photodynamic melting of phase-reversed charge stripes and enhanced condensation

Abstract
The interplay between charge stripes and pairing is a longstanding point of scrutiny in a broad class of unconventional superconductors since, in some cases, it is unclear whether their intertwining benefits the ensuing superfluidity. Experiments that explore the out-of-equilibrium dynamics of these systems try to tip the balance in favor of one phase or the other by selective coupling to relevant modes. Leveraging the fact that competition between stripes and pairing is not exclusive to fermionic systems, we explore the photoirradiation dynamics of interacting hardcore bosons, in which density wave phase-reversal melting gives rise to enhanced superfluid properties, quantified by the dynamic amplification of zero-momentum occupancy and charge stiffness. Our results, obtained using unbiased methods for an interacting system on a ladder geometry, demonstrate how one can engineer time-dependent perturbations to release suppressed orders, potentially providing insight into the underlying mechanism in related experiments.
Introduction.—  Time-dependent experiments provide another angle in the longstanding quest to understand the mechanism binding otherwise repulsive fermions, supporting the emergence of unconventional superconductivity. In particular, rapid energy insertion in the form of an ultrafast laser pump has arisen as a capable technique [1, 2] that coherently drives the dynamics of the electrons in ways that can control magnetic [3, 4], charge [5, 6, 7, 8], and superconducting orders [9, 10, 11, 12, 13, 14, 15]. Owing to the entwining of the different orders that permeate the doped cuprate physics [16, 17], such an approach is one in which coupling to certain degrees of freedom can suppress specific ordering while releasing others [2, 18].
For instance, in equilibrium, the presence of charge stripes, quasi-one-dimensional regions of doped charges (either holes or electrons), is ubiquitous in a large class of cuprate superconductors, often accompanied by phase reversal of local order, such as antiferromagnetic spin alignment [19, 20]. Evidence abounds that when such density modulation is static, it directly hurts superconductivity, as in the case of La2-xBaxCuO4 cuprates [21], in particular at $x=1/8$-doping [22, 23, 24, 25]. Nonetheless, for many other compounds, including bismuth- [26, 27, 28] and yttrium-based cuprates [29, 30, 31, 32], fluctuating stripes associated with incommensurate ordering wave-vectors have been observed, and these often can coexist with coherent pairing.
Although the physics of stripes in equilibrium, including fluctuating/incommensurate ones, is now well studied within a broad class of models and numerical methods [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], direct indication of the dynamical interplay between charge ordering and pairing is nonexistent in unbiased numerics of many-body models even though this is the precise mechanism explored in experiments [9, 43, 11, 44] – we aim to fill this void  [^1] We note microscopic models for cuprates have been studied under photoirradiation [84, 85, 86, 87], but the interplay with stripe physics is currently missing.. Our key observation is that the stripe physics and antiphase of local order are not intrinsic to repulsive interactions [46], nor to fermionic statistics [47], including the fluctuating-static dichotomy [47]. This fundamental, and highly non-trivial, analogy allows one to unbiasedly solve the dynamics of Hamiltonians that contain specific ingredients of actual cuprate physics — namely, strong interactions, superfluidity, and antiphase correlations across stripes — even if the pairing is isotropic, unlike what is observed in a large class of unconventional superconductors.
Our main result is that by adjusting the characteristics of the time-dependent perturbation in the form of a pump in a strongly correlated system, one can show that an effective nonlinear optical coupling is responsible for exciting the system from the ground state in a way that a static stripe with antiphase ordering melts while (quasi-)superfluidity is enhanced dynamically.
Model.—  Our starting point is the Hubbard model for repulsive hardcore bosons,
$\hat{\cal H}=-t_{h}\sum_{\langle{\bf i,j}\rangle}\hat{b}_{\bf i}^{\dagger}\hat{b}_{\bf j}^{\phantom{\dagger}}+V\sum_{\langle{\bf i,j}\rangle}\hat{n}_{\bf i}\hat{n}_{\bf j}+V_{0}\sum_{{\bf i}\in\mathcal{P}}\hat{n}_{\bf i}\ ,$ (1)
where $\hat{b}_{\bf i}^{\phantom{\dagger}}$ ($\hat{b}_{\bf i}^{\dagger}$) annihilates (creates) a hardcore boson at site $\bf i$}, satisfying the statistics $\hat{b}_{\bf i}^{\phantom{\dagger}},\hat{b}_{\bf j}^{\dagger}]=0$ if ${\bf i}\neq{\bf j}$ and $\{\hat{b}_{\bf i}^{\phantom{\dagger}},\hat{b}_{\bf i}^{\dagger}\}=1$ [^2] The hardcore constraint is further imposed by $(\hat{b}_{\bf i}^{\dagger})^{2}=0$ and $(\hat{b}_{\bf i}^{\phantom{\dagger}})^{2}=0$, of an $L_{x}\times L_{y}$ ladder with periodic boundary conditions – we focus on the case $L_{x}=2L_{y}=8$ [Fig. 1(a)]; $\hat{n}_{\bf i}=\hat{b}_{\bf i}^{\dagger}\hat{b}_{\bf i}^{\phantom{\dagger}}$ is the associated number operator. The nearest-neighbor, $\langle{\bf i},{\bf j}\rangle$, hopping energy scale is given by $t_{h}$, and a contact interaction proportional to $V$ establishes the repulsion among the particles. These are supplemented by a modulated chemical potential of amplitude $V_{0}$, constrained to obey $\mod{(i_{x},\mathcal{P})}\equiv 0$, where $\mathcal{P}=4$, as a way to induce static stripe formation explicitly with periodicity $\cal P$} – see Fig. 1(a). While unnecessary in larger ladders [47], it aids in triggering a density modulation characteristic of stripe physics in a relatively small system size.
Figure 1: (a) Schematic illustration of Eq. (1) with relevant terms annotated, subjected to $x$-polarized photoirradiation. Here, the hardcore bosons are represented by a composite local fermion. Momentum distribution profiles $n_{\bf k}$ focusing on $k_{y}=0$ (b) and $k_{y}=\pi$ (c), comparing the equilibrium (before pump) with the out-of-equilibrium (after pump, averaged results from $t\in[11t_{d},26t_{d}]=[72,169]t_{h}^{-1}$; shading gives the standard deviation) response to the external field. We observe an average relative enhancement of about 37% for the zero-momentum condensation under the photoirradiation. (d) The $x$-direction charge stiffness resolved for the first 200 eigenstates $|m\rangle$ of $\hat{\cal H}$; marker colors are mapped by the overlap of the time evolved state at $t_{1}=5t_{d}$ and the corresponding $|m\rangle$. (e) The instantaneous energy $E(t)=\langle\psi(t)|\hat{\cal H}(t)|\psi(t)\rangle$ over the dynamics; the shaded region marks the probing for the dynamical charge stiffness starting at $t_{1}$, with a linear in time vector potential with slope $\delta A_{x}$ – top panel depicts the total ${\bf A}(t)$. (f) Energy variation in the probe regime as a function of the effective flux $\Phi_{x}(t)=\delta A_{x}L_{x}(t-t_{1})$ for $t>t_{1}$$\delta A_{x}=0.01$. Dashed line displays an approximant at small fluxes that quantifies the dynamical charge stiffness at $t=t_{1}$, confirming a robust charge transport induced by the pump (see text). Pulse parameters are $\Omega=1.713\ t_{h}$, $A_{0}=0.62$ and $t_{d}=6.5t_{h}^{-1}$.
The dynamics under photoirradiation are emulated by including a time-dependent vector potential, ${\bf A}(t)$. As with fermions, hardcore bosons couple to ${\bf A}(t)$, provided they have an associated charge, through the Peierls’ substitution, $\hat{b}^{\dagger}_{{\bf i}}\hat{b}^{\phantom{\dagger}}_{{\bf j}}\rightarrow e^{i\mathbf{A}(t)\cdot(\mathbf{i}-\mathbf{j})}\,\hat{b}^{\dagger}_{{\bf i}}\hat{b}^{\phantom{\dagger}}_{{\bf j}}$. Focusing on ultrafast pumps, the light pulse is modeled as ${\bf A}(t)=A_{0}e^{-t^{2}/2t_{d}^{2}}\cos(\Omega t+\phi_{t}){\bf e}_{\rm pol}$; assigning a temporal width $t_{d}$, amplitude $A_{0}$, frequency $\Omega$, and $\phi_{t}$ is a temporal phase shift. The polarization direction is chosen as $^{\bf e}_{\mathrm{pol}}=\hat{\bf x}$, i.e., perpendicular to the stripes — see Fig. 1(a). The unitary dynamics, $|\psi(t+dt)\rangle=e^{-{i}\hat{\mathcal{H}}(t)dt}|\psi(t)\rangle$ $\hbar\equiv 1$], is computed via Krylov subspace methods [49, 50], by taking $|\psi(t\to-\infty)\rangle=|0\rangle$, the equilibrium ground-state of (1) in the zero quasi-momentum sector where it resides for our set of parameters  [^3] Since the pump does not break translation invariance, this is the sector the dynamics explore..
Enhanced condensation and dissipationless transport.—  Interacting hardcore bosons at $T=0$ give rise to either charge ordered phases or a superfluid phase [52, 53, 54, 55, 56, 57], depending on the interaction-hopping ratio $V/t_{h}$ — in what follows we set $V/t_{h}=4$. Tendency towards phase separation and stripe formation in the vicinity of half-filling, $\langle\hat{n}\rangle=1/2$, has been reported [55, 57], and stabilization of such a regime can be made explicit via an external periodic potential [47] in the doped regime  [^4] Including an extra potential has been used in other contexts to stabilize magnetic ordering for example [88]. Notably, in this scenario, the propensity for spatial organization of holes is accompanied by a reversal in the phase of the checkerboard density modulation across a stripe boundary provided the corresponding chemical potential $V_{0}$ is sufficiently large – this is often dubbed as a $\pi$-phase shift [33, 46, 47]. Focusing on the hole density $\delta\equiv 1/2-\langle\hat{n}\rangle=1/16$, Fig. 2(a) displays the connected density-density correlations across a stripe, $c_{{\bf i}_{0}}(dx,dy)=\langle\hat{n}_{{\bf i}_{0}}\hat{n}_{{\bf i}+d{\bf r}}\rangle-\langle\hat{n}_{{\bf i}_{0}}\rangle\langle\hat{n}_{{\bf i}+d{\bf r}}\rangle$; ${\bf i}_{0}$ is a reference site and $d{\bf r}\equiv(dx,dy)$ the spatial displacement. As $V_{0}/t_{h}$ increases, the same sublattice correlation $c_{{\bf i}_{0}}(2,0)$ across a stripe transitions from positive to negative row whereas $c_{{\bf i}_{0}}(2,1)$ shows the opposite trend, establishing a direct verification of the antiphase charge ordered pattern once hole stripes are sufficiently robust.
At the same time, a measure of (quasi-)condensation is given by the momentum distribution function $n_{{\bf k}}=\frac{1}{L_{x}L_{y}}\sum_{{\bf i},{\bf j}}e^{-i{\bf k}\cdot({\bf i}-{\bf j})}\langle\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}\rangle$, which, at zero temperature, exhibits dependence that is extensive in the system size for the zero-momentum occupancy, $n_{{\bf k}=0}$. Figure 2(a) shows that $n_{{\bf k}=0}$ is suppressed with the emergence of a density wave phase reversal in the ground state $|0\rangle$ at large $V_{0}$’s. Such anticorrelation is present in reverse across selected state in the energy spectrum, particularly for the second excited state $|2\rangle$ [Fig. 2(b)]  [^5] The eigenstates listed here obey translation invariance with periodicity ${\cal P}=4$ in $x$-direction and 1 in the $y$-direction and reside in the same quasi-momentum ${\bf q}=0$.: the $\pi$-phase shift is absent and $n_{{\bf k}=0}$ is relatively robust, especially at large $V_{0}/t_{h}$ [we focus on $V_{0}=5t_{h}$ in what follows]. This suggests that fine-tuned perturbations tailored to access the physics of this (and potentially other) excited state(s) are good starting points to tip the balance in favor of enhanced superfluidity. Figures 1(b) and 1(c) report our main results: The zero-momentum occupancy can be made substantially larger ($\simeq 37\%$) after photoirradiation, at the expense of reduction of the occupation of other momentum modes [Fig. 1(c)].
Figure 2: Dependence of the density-density correlations across a stripe for the ground state (a) and the second excited state (b) on the stripe potential. The solid vertical line indicates our selected parameter set for the dynamics, $V_{0}/t_{h}=5$, featuring a robust charge $\pi$-phase shift for the ground state, which is absent in the excited state. The right vertical axes display the zero-momentum occupancies.
Nonetheless, condensation need not be necessarily tied to superfluid behavior, in general [60]. A necessary condition for dissipationless charge transport  [^6] But not sufficient for a finite superfluidity, which entails phase coherence, quantified by a finite superfluid weight. is given by the charge stiffness (or Drude weight), whose finite value indicates an ideal metal with ballistic transport [62, 63, 64]. For the $m$-th eigenstate of $\hat{\cal H}$ with eigenenergy $E_{m}$,
$\frac{D_{m}^{(\mu)}}{\pi}=\frac{L_{\mu}^{2}}{L_{x}L_{y}}\left.\frac{\partial^{2}E_{m}}{\partial\Phi_{\mu}^{2}}\right|_{\Phi_{\mu}=0}\ ,$ (2)
where $\mu=x,y$ and $\Phi_{\mu}=\int_{0}^{L_{\mu}}{\bf A}\cdot d{\bf l}_{\mu}$ the corresponding flux associated with the $\mu$-direction for a given system size. At finite-temperature $T$ [65, 66] or under out-of-equilibrium [67] conditions, $D=\sum_{m}p_{m}D_{m}$, with $p_{m}=e^{-E_{m}/T}/{\cal Z}$, the Boltzmann weight in the former $k_{B}\equiv 1$], and $p_{m}=|\langle m|\psi(t)\rangle|^{2}$ in the latter. Figure 1(d) shows the eigenstate resolved charge stiffness longitudinal to the ladder using a numerical differentiation of the eigenenergies $E_{m}$ with respect to $\Phi_{x}$. That the roughly equal distribution of $D_{m}^{(x)}$ about zero is observed across the spectrum is an indication of the vanishing thermal averaged charge stiffness, as similarly indicated in the one-dimensional fermionic Hubbard model [68, 69, 70, 71].
Under out-of-equilibrium conditions, quantification of charge transport can be accomplished by threading a dynamical flux $\Phi_{x}(\Delta t)=\delta A_{x}L_{x}\Delta t$, associated with an electric field $\frac{\partial(\Phi_{x}(t)/L_{x})}{\partial t}=\delta A_{x}$ along the $x$-direction during a time-interval $\Delta t$ [67]. Such a linear-in-time probing vector potential, $^{\bf A}(t)=\delta A_{x}(t-t_{1})\hat{x}$ for $\Delta t=t-t_{1}>0$, is turned on, in our simulations, much after the center of the photoirradiation pulse, i.e., $t_{1}\gg t_{d}$ [see a representation on Fig. 1(e), top], to quantify if the pulse endowed dissipationless transport properties. This can be obtained in two ways: The first is to recognize that the instantaneous energy $E(t)\equiv\langle\psi(t)|\hat{\cal H}(t)|\psi(t)\rangle$ [Fig. 1(e)] can, at short-times (or probing fluxes), be expanded as $E(t)=E(t_{1}+\Delta t)\simeq E(t_{1})+\frac{1}{2}\left.\frac{\partial^{2}E(t_{1}+\Delta t)}{\partial\Phi_{x}^{2}}\right|_{\Delta t\to 0}\Phi^{2}_{x}(\Delta t)$. As such, the energy gain $\Delta E=E(t_{1}+\Delta t)-E(t_{1})$ associated with the probing flux is $\Delta E=\frac{D^{(x)}}{4\pi}\Phi_{x}^{2}(\Delta t)$, where we have used the definition of the Drude weight [Eq. (2)] for energy $E$ and $L_{x}=2L_{y}$.
Figure 1(f) shows the time-dependence of the energy gain during the probe time. Indeed, at short times, a quadratic dependence in $\Phi_{x}(\Delta t)$ emerges, indicating a finite $D^{(x)}$ and thus dissipationless transport after the pulse. The second estimation of this physical behavior is that $D^{(x)}$ closely matches with an approximant of the out-of-equilibrium Drude weight $\tilde{D}^{(x)}=\sum_{m=0}^{199}|\langle m|\psi(t_{1})\rangle|^{2}D_{m}^{(x)}=0.180$, by taking the lowest 200 lowest energy eigenstates of the equilibrium Hamiltonian, which concentrates most of the weight of $|\psi(t)\rangle$ at the time $t=t_{1}$ of the application of the probing flux — overlaps $|\langle m|\psi(t_{1})\rangle|^{2}$ are color-mapped in Fig. 1(d). This is to be contrasted against the equilibrium Drude weight, $D_{0}^{(x)}=0.026$, an order-of-magnitude smaller, furnishing evidence for photoinduced ballistic transport for this system size.
In addition, the energy oscillations in the probing regime with periodicity $2\pi$ in the applied flux [see vertical dotted lines in Fig. 1(f)] confirm that the fundamental charge in these conditions is $q=1$, a single hard-core boson. This is a direct generalization to the dynamical setting of the Byers-Yang theorem [72, 73], establishing that the energy of a system of particles with charge $q$ confined to a closed geometry and subjected to an external flux exhibits periodicity in $\Phi$ with period $\Phi_{0}=2\pi/q$, the flux quantum.
Figure 3: (a) Gaussian pump field with amplitude $A_{0}=0.62$, frequency $\Omega=\omega_{26}\simeq 1.713\ t_{h}$, and temporal width $t_{d}=6.5t_{h}^{-1}$ — the same as in Fig. 1. Phase-averaged dynamics (solid curves) for a range of phases $\phi_{t}\in[0,2\pi)$ (shaded regions) are shown in (b)–(d). (b) Dynamics of the correlation functions $c_{{\bf i}_{0}}(2,0)$ and $c_{{\bf i}_{0}}(2,1)$, in respect to the reference site ${\bf i}_{0}$ marked in Fig. 1(a), exhibiting $\pi$-phase shift reversal after the pulse. (c) The zero-momentum occupation $n_{{\bf k}=0}$ exhibits a 37% enhancement in comparison to the equilibrium. (d) State projection coefficients for selected equilibrium eigenstates $|m\rangle$ (the ones showing significant weight over the dynamics), with state $|2\rangle$ gradually becoming dominant at long times, added by an enhanced overlap with state $|6\rangle$.
Melting of $\pi$-phase shift.—  Having confirmed that an appropriate photoirradiation leads to not only a larger occupancy of the zero-momentum mode but also enhanced transport, we now focus on the fate of the charge $\pi$-shift mediated by the stripes. Under the same photoirradiation conditions, later detailed, and apart from the introduction of a probe vector potential, Figure 3 shows the dynamics of $n_{{\bf k}=0}$ closely — data are reported for a range of the pump’s phase shifts $\phi_{t}$, as well as their average, which aids in smoothening the short-time coherent oscillations. That the photoirradiation can be used to filter pairing in detriment of charge-stripes is apparent by the fact that while $n_{{\bf k}=0}$ dynamically increases [Fig. 3(c)], the $\pi$-phase shift disappears: $c_{{\bf i}_{0}}(2,0)$ and $c_{{\bf i}_{0}}(2,1)$ switch their signs [Fig. 3(b)] at the time at which the pump [Fig. 3(a)] has the largest amplitude.
This can be understood in terms of the overlaps $|\langle m|\psi(t)\rangle|^{2}$ with low-lying equilibrium eigenstates $|m\rangle$ in Fig. 3(d). At long times, the dynamics stabilize most of the spectral weight in state $|2\rangle$, the target state, as well as $|6\rangle$ with a still relevant weight ($13\%$) with the initial state, $|0\rangle$. For that, one needs to optimize pump parameters, including the pump’s width $t_{d}$ and amplitude $A_{0}$, as indicated in Figs. 4(a) and 4(b), resulting in optimal $A_{0}=0.62$, and $t_{d}=6.5t_{h}^{-1}$. However, additional setting of the pump’s frequency is required to achieve the most accurate targeting of the excited state. Previous investigations in various models [74, 75, 76] established that a resonant photoirradiation, $\Omega=E_{\rm target}-E_{0}$, could accomplish a direct targeting of relevant excited states with eigenenergies $E_{\rm target}$. At the crux of such a process is the fact that first-order time-dependent perturbation theory stemming from the pulse application (details in the Supplemental Materials [77]) establishes that the $x$-direction current operator, $\hat{J}_{x}=it_{h}\sum_{\bf i}(\hat{b}_{\bf i+\hat{x}}^{\dagger}\hat{b}_{\bf i}^{\phantom{\dagger}}-b_{\bf i}^{\dagger}\hat{b}_{\bf i+\hat{x}}^{\phantom{\dagger}})$, connects these states with a robust $|\langle m_{\rm target}|\hat{J}_{x}|0\rangle|$. Unfortunately, our investigation shows that, in our case, the matrix elements of the current operator between $|0\rangle$ and $|2\rangle$ are small [Fig. 4(c)].
Figure 4: Dependence on the pump width $t_{d}$ and amplitude $A_{0}$ of the long-time overlap at $t_{\rm max}=9t_{d}$ with selected equilibrium eigenstates, $|m=2\rangle$ (a) and $|m=6\rangle$ (b). A finer mesh resolution is shown in the area with an enhanced overlap $|\langle\psi(t_{\rm max})|2\rangle|^{2}$ while the cross marker pinpoints the maximum value. Here, the pump frequency is set at $\Omega=\omega_{26}=1.713\ t_{h}$. (c) The equilibrium $x$-direction current matrix in the low-lying spectrum; dashed lines highlight matrix elements with the target state $|2\rangle$.
The non-linear optical coupling.—  That resonant targeting cannot couple to the target state does not impede higher-order, and thus nonlinear, optical excitation from taking place. This is particularly possible because intermediate equilibrium states can have a relatively large current matrix element, e.g., $|\langle 0|\hat{J}_{x}|6\rangle|\simeq 0.66$ and $|\langle 6|\hat{J}_{x}|2\rangle|\simeq 0.58$. Indeed, in second-order time-dependent perturbation theory, the time-dependent expansion coefficients in the interaction picture entering in $|\psi_{\rm I}(t)\rangle=\sum_{m}c_{m}(t)|m\rangle$, for initial state $|0\rangle$, final state $|2\rangle$, intermediate state $|m_{1}\rangle$, at infinite time, are (see Ref. [77])
$c_{m=2}^{(2)}(\infty)=\sum_{m_{1}}\sum_{n_{1},n_{2}}$ $\,\hat{V}_{2,m_{1}}^{(n_{2})}\hat{V}_{m_{1},0}^{(n_{1})}\cdot\frac{1}{E_{m_{1}}-E_{0}-n_{1}\Omega-i\epsilon}$
$\times\delta\left(E_{2}-E_{0}-(n_{1}+n_{2})\Omega\right)\ ,$ (3)
where $\hat{V}_{m,m^{\prime}}^{(n)}$ are the matrix elements of the $n$-th Fourier component of the perturbation in the unperturbed eigenstates $|m\rangle$ of $\hat{\cal H}$, and the delta function enforces total energy conservation via the absorption or emission of $n_{1}+n_{2}$ photons. Now, considering the intermediate state $|m_{1}=6\rangle$ has a significant contribution, a close resonance occurs for $E_{6}\approx E_{0}+n_{1}\Omega$, when the denominator in Eq. (3) is small. Meanwhile, the energy conservation condition yields $E_{2}=E_{0}+(n_{1}+n_{2})\Omega$. Figure 5(a) shows that at pump frequencies $\Omega=\omega_{62}\equiv E_{6}-E_{2}$, a large overlap with the target state $|2\rangle$ is obtained at long times. Using the two previous energy conditions, $\omega_{62}=-n_{2}\Omega$, indicating that $n_{2}=-1$ (a photon emission in between the intermediate and final states), explains the observations. As a result, the resonance condition leads to $E_{6}\approx E_{0}+n_{1}\omega_{62}$, therefore, $\omega_{60}/\omega_{62}\approx n_{1}$. The equilibrium eigenspectrum yields $\omega_{60}/\omega_{62}\simeq 2.75$, indicating that the resonant excitation process from the ground-state to the intermediate one likely involves the absorption of three quanta of energy $\Omega=\omega_{62}$.
Figure 5: Dependence on the pump frequency $\Omega$ and amplitude $A_{0}$ of the long-time overlap at $t_{\rm max}=9t_{d}=27t_{h}^{-1}$ with selected equilibrium eigenstates, $|m=2\rangle$ (a), $|m=3\rangle$ (b), and $|m=6\rangle$ (c). Horizontal lines indicate relevant gaps in the equilibrium spectrum; note the different ranges of the color bars, which are set to enhance visualization.
Lastly, to confirm this perturbation picture, the perturbation matrix elements explain an enlarged overlap at long times with other states as well, but at first-order: a considerable overlap with state $|6\rangle$ [Fig. 5(c)] is obtained at the resonance $\omega_{60}/\Omega_{\rm pump}=n_{1}=4.69/1.57\simeq 3$, via a three quanta absorption. In turn, the most prominent matrix element of $\hat{J}_{x}$ in the low-lying spectrum, $|\langle 3|\hat{J}_{x}|0\rangle|\simeq 1.93$, can be dynamically accessed in first-order with pump frequencies $\Omega_{\rm pump}\simeq\omega_{30}$ — Figure 5(b) precisely confirms this scenario.
Summary and outlook.—  We showed that fine-tuned photoirradiation, effectively promoting a non-linear optical excitation towards a target state, forms a pathway for dynamically tailoring a system to exhibit desired properties. This was exemplified in the case of the competing interplay between pairing and density-wave with phase reversal in the presence of stripes; however, the approach is likely applicable in other scenarios. Key is the identification of the equilibrium properties of the low-lying eigenspectrum, and the design of a perturbation that, in sufficiently low order, can access them. Notably, that generically melting stripes leads to transient superconductivity is a result under intense debate in experiments [78]. Our results affirmatively point to this connection in a specific model. Confirming this for the fermionic Hubbard and $t$-$J$ models relevant to the cuprates, and larger system sizes, requires numerics currently beyond the capabilities of exact diagonalization and Krylov methods we employ. This is currently only possible with other techniques that approximate the dynamics, such as the time-dependent variational principle in the language of matrix product states [79, 80, 81, 82] — we leave this for future investigation.
Acknowledgements.
R.M. acknowledges support from the TcSUH Welch Professorship Award. R.T.S. is supported by the grant DOE DE-SC0014671 funded by the U.S. Department of Energy, Office of Science. Numerical simulations were performed in the Tianhe-2JK at the Beijing Computational Science Research Center and with resources provided by the Research Computing Data Core at the University of Houston. This work also used TAMU ACES at Texas A&M HPRC through allocation PHY240046 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296. The data that support the findings of this article are openly available [83].
Supplementary Materials
Photodynamic melting of phase-reversed charge stripes and enhanced condensation
Supplementary Materials
Photodynamic melting of phase-reversed charge stripes and enhanced condensation
These supplementary materials present additional results, including a detailed perturbative analysis of the photoirradiation, which justifies the investigation of matrix elements of current and kinetic energy operators. Analysis of such matrix elements also corroborates the choice of $x$-polarization, i.e., polarization perpendicular to the stripe directions. Additional investigations include the time evolution of the density during photodynamics and the usage of a pump resonant with other intermediate states.
1 Time-dependent Perturbation Theory for the photoirradiation
The main text explains the enhancement of zero-momentum occupancy and the melting of the charge $\pi$-phase shift by targeting eigenstates of the low-lying equilibrium energy spectrum. The manner by which this targeting is achieved is through the tuning of the pulse properties, in particular its frequency, such that at long times, the weight of the target state in the time-evolving wavefunction $|\psi(t)\rangle$ is maximized. We now explain how this arises as a consequence of time-dependent perturbation theory applied to the pulsed driving — Ref. [90] describes an equivalent analysis in another lattice model.
The effect of the pulse is to modify the equilibrium Hamiltonian $\hat{\cal H}$ [Eq. (1) of the main text] with a time-dependent perturbation $\hat{V}(t)$,
$\hat{H}(t)=\hat{\cal H}+\hat{V}(t)\ .$ (S1)
Considering that the eigenstates $|m\rangle$ of $\hat{\cal H}$ define a basis, the time-evolved state at arbitrary times takes the form
$|\psi(t)\rangle=\sum_{m}c_{m}(t)e^{-iE_{m}t}|m\rangle\ ,$ (S2)
where $c_{m}(t)$ are the time-dependent expansion coefficients in the Schrödinger picture, related to the interaction picture amplitudes by $\langle m|\psi_{\rm I}(t)\rangle=e^{iE_{m}t}c_{m}(t)$. In this latter picture, the state vector evolves according to,
$i\frac{d}{dt}|\psi_{\rm I}(t)\rangle=\hat{V}_{\rm I}(t)|\psi_{\rm I}(t)\rangle,$ (S3)
with $\hat{V}_{\rm I}(t)=e^{i\hat{\cal H}t}\hat{V}(t)e^{-i\hat{\cal H}t}$ and $|\psi_{\rm I}(t)\rangle=e^{i\hat{\cal H}t}|\psi(t)\rangle$. In particular, recalling that the ground-state $|0\rangle$ coincides with the initial state of the system, $|\psi(t\to-\infty)\rangle$, the interaction picture state can be expanded in a Dyson series [91]:
$|\psi_{\rm I}(t)\rangle$ $=\mathcal{T}\exp\left(-i\int_{-\infty}^{t}dt^{\prime}\hat{V}_{\rm I}(t^{\prime})\right)|0\rangle$
$=|0\rangle+|\psi_{\rm I}^{(1)}(t)\rangle+|\psi_{\rm I}^{(2)}(t)\rangle+\ldots\ ,$ (S4)
where $\mathcal{T}$ is the time-ordering operator. To first order in the perturbation, one obtains
$|\psi_{\rm I}^{(1)}(t)\rangle=-i\int_{-\infty}^{t}dt_{1}\,\hat{V}_{\rm I}(t_{1})|0\rangle\ ,$ (S5)
whereas the second-order contribution yields,
$|\psi_{\rm I}^{(2)}(t)\rangle=(-i)^{2}\int_{-\infty}^{t}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,\hat{V}_{\rm I}(t_{2})\hat{V}_{\rm I}(t_{1})|0\rangle\ .$ (S6)
Focusing on the first order, the amplitude to find the system in the unperturbed state $|m\rangle$, $\langle m|\psi_{\rm I}^{(1)}(t)\rangle\equiv c_{m}^{(1)}(t)$, becomes
$c_{m}^{(1)}(t)=-i\int_{-\infty}^{t}dt_{1}\,e^{iE_{m}t_{1}}\langle m|\hat{V}(t_{1})|0\rangle e^{-iE_{0}t_{1}}\ .$ (S7)
In turn, the second-order amplitude reads,
$c_{m}^{(2)}(t)=(-i)^{2}\sum_{m^{\prime}}$ $\int_{-\infty}^{t}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,$
$\times e^{iE_{m}t_{2}}\langle m|\hat{V}(t_{2})|m^{\prime}\rangle e^{-iE_{m^{\prime}}t_{2}}$
$\times e^{iE_{m^{\prime}}t_{1}}\langle m^{\prime}|\hat{V}(t_{1})|0\rangle e^{-iE_{0}t_{1}}\ ,$ (S8)
where the index $m^{\prime}$ runs over all intermediate eigenstates of the unperturbed Hamiltonian.
If the perturbation is rewritten in terms of its Fourier series,
$\hat{V}(t)=\sum_{n=-\infty}^{\infty}\hat{V}_{n}e^{-in\Omega t},$ (S9)
then, in the infinite-time limit, the first-order transition amplitude [Eq.(S7)], after the time integration  [^7] Using the identity $\int_{-\infty}^{\infty}dte^{i\omega t}=2\pi\delta(\omega)$, becomes
$c_{m}^{(1)}(\infty)=-2\pi i\sum_{n}\langle m|\hat{V}_{n}|0\rangle\,\delta(E_{m}-E_{0}-n\Omega)\ ,$ (S10)
which reflects energy conservation for the absorption and emission of $n$ photons of frequency $\Omega$.
Similarly, the second-order transition amplitude becomes
$c_{m}^{(2)}(\infty)=-\sum_{n_{1},n_{2}}\sum_{m^{\prime}}$ $\frac{\langle m|\hat{V}_{n_{2}}|m^{\prime}\rangle\langle m^{\prime}|\hat{V}_{n_{1}}|0\rangle}{E_{m^{\prime}}-E_{0}-n_{1}\Omega-i\epsilon}\,$
$\times\delta(E_{m}-E_{0}-(n_{1}+n_{2})\Omega)\ ,$ (S11)
where the sum runs over all intermediate equilibrium states $|m^{\prime}\rangle$, and the delta function enforces energy conservation for the net exchange of $n_{1}+n_{2}$ photons.
Now that we have established the perturbative coefficients, we are in a position to describe their complete functional form. First, we recall the structure of $\hat{V}(t)$ in Eq. (S1), which arises from the hopping modification of the kinetic part of the Hamiltonian $\hat{H}(t)=\hat{K}(t)+\hat{V}(t)$]:
$\hat{K}(t)$ $=-t_{h}\sum_{\langle{\bf i},{\bf j}\rangle}\Big{[}e^{iA_{{\bf ij}}(t)}\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}+e^{-iA_{{\bf ij}}(t)}\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\Big{]}$
$=-t_{h}\sum_{\langle{\bf i},{\bf j}\rangle}\Big{[}\cos A_{{\bf ij}}(t)\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$\quad\quad\quad\quad+i\sin A_{{\bf ij}}(t)\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\Big{]}\ ,$ (S12)
where $A_{{\bf ij}}(t)={\bf A}(t)\cdot({\bf i}-{\bf j})$. As a result, the time-dependent perturbation in Eq. (S1) yields,
$\hat{V}(t)=-t_{h}\sum_{\langle{\bf i},{\bf j}\rangle}\Big{[}\left(\cos A_{{\bf ij}}(t)-1\right)$ $\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$+i\sin A_{{\bf ij}}(t)$ $\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{\bf j}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{\bf j}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\Big{]}\ ,$ (S13)
which is a sum of a time-modulated kinetic-energy and current-like terms for the direction in which the vector potential is applied. For instance, for our pulse polarized along the $x$-direction, $^{\bf A}(t)=A(t)\hat{\bf x}$, $A_{{\bf ij}}(t)=A(t)$ if $^{\bf j}={\bf i}+\hat{\bf x}$, and $A_{{\bf ij}}(t)=0$ if $^{\bf j}={\bf i}+\hat{\bf y}$. As a result, the time-dependent perturbation can be written as $\hat{V}(t)=\hat{K}_{x}(t)+\hat{J}_{x}(t)$, where
$\hat{K}_{x}(t)\equiv-t_{h}\left(\cos A(t)-1\right)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$ (S14)
and,
$\hat{J}_{x}(t)\equiv-it_{h}\sin A(t)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\ .$ (S15)
Now there are two choices to achieve analytical expressions for the transition amplitudes: The first is to assume the perturbation is perfectly time-periodic, using the discrete Fourier transform [Eq. (S9)] to fully describe the Floquet-like driving of the system, i.e., by taking the pulse width $t_{d}\to\infty$, i.e., $A(t)\simeq A_{0}\cos(\Omega t)$ and the second is to taking into account the Gaussian envelope that modulates the sinusoidal pulse function. In the former fully periodic case, the discrete Fourier coefficients in Eq. S9, read
$\hat{V}_{n}=\frac{1}{T}\int_{0}^{T}dt\ \hat{V}(t)e^{in\Omega t},\quad\text{with,}\quad T=2\pi/\Omega.$ (S16)
Equivalently, one can deduce these coefficients by making use of the Jacobi-Anger expansions [89], for the relevant terms in $\hat{K}_{x}(t)$ and $\hat{J}_{x}(t)$:
$\cos(A_{0}\cos(\Omega t))-1$ $=\mathcal{J}_{0}(A_{0})-1$
$\quad+2\sum_{n=1}^{\infty}\mathcal{J}_{2n}(A_{0})\cos(2n\Omega t)$
$=\sum_{m=-\infty}^{\infty}a_{m}e^{-im\Omega t}\ ,$ (S17)
and
$\sin(A_{0}\cos(\Omega t))$ $=2\sum_{n=0}^{\infty}\mathcal{J}_{2n+1}(A_{0})\sin((2n+1)\Omega t)$
$=\sum_{m=-\infty}^{\infty}b_{m}e^{-im\Omega t}\ ,$ (S18)
where $\mathcal{J}_{n}(x)$ are the Bessel functions of the first kind of integer order $n$. After expanding $\cos(2n\Omega t)$ and $\sin((2n+1)\Omega t)$ to their exponential form, it leads to the identification $a_{0}={\cal J}_{0}(A_{0})-1$, $a_{\pm 2n}={\cal J}_{2n}(A_{0})$, $a_{\pm(2n+1)}=0$ and $b_{\pm(2n+1)}=\mp i{\cal J}_{2n+1}(A_{0})$, $b_{2n}=0$. This fully defines thus the terms entering in the transition amplitudes in Eqs. (S10) and (S12) as,
$\hat{V}_{n}=$ $-t_{h}a_{n}\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$-it_{h}b_{n}\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)\ .$ (S19)
In the second case, where $A(t)=A_{0}e^{-t^{2}/2t_{d}^{2}}\cos(\Omega t)$, instead of using the Fourier expansion  [^8] Which for this case should be instead a continuous Fourier transform unlike the discrete case of Eq. (S9)., insight can be gained by taking the limit of small pulse amplitudes, $A_{0}\ll 1$, such that $\cos A(t)\approx 1-\frac{1}{2}A^{2}(t)$ and $\sin A(t)\approx A(t)$, leading to
$\hat{K}_{x}(t)$ $\approx\frac{t_{h}}{2}A^{2}(t)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$=\frac{t_{h}}{2}A_{0}^{2}e^{-t^{2}/t_{d}^{2}}[1+\cos(2\Omega t)]\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}+\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$\equiv A^{2}(t)\hat{K}_{x}^{(2)}\ ,$ (S20)
and
$\hat{J}_{x}(t)$ $\approx-it_{h}A(t)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$=-it_{h}A_{0}e^{-t^{2}/2t_{d}^{2}}\cos(\Omega t)\sum_{\bf i}\left(\hat{b}^{\dagger}_{\bf i}\hat{b}_{{\bf i}+\hat{\bf x}}^{\phantom{\dagger}}-\hat{b}^{\dagger}_{{\bf i}+\hat{\bf x}}\hat{b}_{\bf i}^{\phantom{\dagger}}\right)$
$\equiv A(t)\hat{J}_{x}^{(2)}\ .$ (S21)
As a result, the first-order transition amplitude [Eq. (S7) for $t\to\infty$] becomes:
$c_{m}^{(1)}(\infty)$ $\simeq-i\int_{-\infty}^{\infty}dt\,\Big{[}A(t)\langle m|\hat{J}_{x}^{(2)}|0\rangle+A^{2}(t)\langle m|\hat{K}_{x}^{(2)}|0\rangle\Big{]}e^{i\omega_{m0}t}$
$=-it_{d}\Bigg{\{}\frac{1}{2}A_{0}\sqrt{2\pi}\left[e^{-\frac{t_{d}^{2}}{2}(\omega_{m0}-\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{2}(\omega_{m0}+\Omega)^{2}}\right]\langle m|\hat{J}_{x}^{(2)}|0\rangle$
$\quad+A_{0}^{2}\frac{\sqrt{\pi}}{2}\left[e^{-\frac{t_{d}^{2}}{4}\omega_{m0}^{2}}+\frac{1}{2}\left(e^{-\frac{t_{d}^{2}}{4}(\omega_{m0}-2\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{4}(\omega_{m0}+2\Omega)^{2}}\right)\right]\langle m|\hat{K}_{x}^{(2)}|0\rangle\Bigg{\}}\ ,$ (S22)
where we use the main text notation $\omega_{m0}\equiv E_{m}-E_{0}$. The resulting transition probability is then governed by the spectral content of $A(t)$ and $A^{2}(t)$, which are sharply peaked around $\Omega$ and $2\Omega$, respectively.
Likewise, in the second-order expansion under the small-$A_{0}$ approximation, the transition amplitude becomes
$c_{m}^{(2)}(\infty)\simeq$ $-\sum_{m^{\prime}}\int_{-\infty}^{\infty}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,e^{i\omega_{mm^{\prime}}t_{2}}e^{i\omega_{m^{\prime}0}t_{1}}$
$\times\Big{[}A(t_{2})A(t_{1})\langle m|\hat{J}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{J}_{x}^{(2)}|0\rangle+A(t_{2})A^{2}(t_{1})\langle m|\hat{J}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{K}_{x}^{(2)}|0\rangle$
$\quad+A^{2}(t_{2})A(t_{1})\langle m|\hat{K}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{J}_{x}^{(2)}|0\rangle+A^{2}(t_{2})A^{2}(t_{1})\langle m|\hat{K}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{K}_{x}^{(2)}|0\rangle\Big{]}$
$=$ $-\sum_{m^{\prime}}\Big{[}\langle m|\hat{J}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{J}_{x}^{(2)}|0\rangle\,\mathcal{I}_{JJ}^{(m^{\prime})}+\langle m|\hat{K}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{K}_{x}^{(2)}|0\rangle\,\mathcal{I}_{KK}^{(m^{\prime})}$
$\quad+\left(\langle m|\hat{J}_{x}^{(2)}|m^{\prime}\rangle\langle m^{\prime}|\hat{K}_{x}^{(2)}|0\rangle+\text{H.c.}\right)\,\mathcal{I}_{JK}^{(m^{\prime})}\Big{]}\ ,$ (S23)
where the time integrals are weighted convolutions of Gaussian wavepackets centered at those harmonics, and are approximated by
$\mathcal{I}_{JJ}^{(m^{\prime})}$ $=\int_{-\infty}^{\infty}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,e^{-\frac{t_{1}^{2}+t_{2}^{2}}{2t_{d}^{2}}}\cos(\Omega t_{1})\cos(\Omega t_{2})\,e^{i\omega_{m^{\prime}0}t_{1}}e^{i\omega_{mm^{\prime}}t_{2}}$
$\approx\frac{\pi t_{d}^{2}}{8}\left[e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}+\Omega)^{2}+(\omega_{mm^{\prime}}+\Omega)^{2}\right]}\right.+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}-\Omega)^{2}+(\omega_{mm^{\prime}}+\Omega)^{2}\right]}$
$\quad+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}+\Omega)^{2}+(\omega_{mm^{\prime}}-\Omega)^{2}\right]}\left.+e^{-\frac{t_{d}^{2}}{2}\left[(\omega_{m^{\prime}0}-\Omega)^{2}+(\omega_{mm^{\prime}}-\Omega)^{2}\right]}\right]\ ,$ (S24)
where we have assumed in the integration that $t_{d}\Omega\gg 1$. Likewise,
$\mathcal{I}_{KK}^{(m^{\prime})}$ $=\int_{-\infty}^{\infty}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,e^{-\frac{t_{1}^{2}+t_{2}^{2}}{t_{d}^{2}}}[1+\cos(2\Omega t_{1})][1+\cos(2\Omega t_{2})]\,e^{i\omega_{m^{\prime}0}t_{1}}e^{i\omega_{mm^{\prime}}t_{2}}$
$\approx\frac{\pi t_{d}^{2}}{2}e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}^{2}+\omega_{mm^{\prime}}^{2})}+\frac{\pi t_{d}^{2}}{4}\left[e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}+2\Omega)^{2}+\omega_{mm^{\prime}}^{2}]}+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+\omega_{mm^{\prime}}^{2}]}\right]$
$+\frac{\pi t_{d}^{2}}{4}\left[e^{-\frac{t_{d}^{2}}{4}[(\omega_{mm^{\prime}}+2\Omega)^{2}+\omega_{m^{\prime}0}^{2}]}+e^{-\frac{t_{d}^{2}}{4}[(\omega_{mm^{\prime}}-2\Omega)^{2}+\omega_{m^{\prime}0}^{2}]}\right]$
$+\frac{\pi t_{d}^{2}}{8}\biggl{[}e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}+2\Omega)^{2}+(\omega_{mm^{\prime}}+2\Omega)^{2}]}+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}+2\Omega)^{2}+(\omega_{mm^{\prime}}-2\Omega)^{2}]}$
$\quad\quad+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+(\omega_{mm^{\prime}}+2\Omega)^{2}]}+e^{-\frac{t_{d}^{2}}{4}[(\omega_{m^{\prime}0}-2\Omega)^{2}+(\omega_{mm^{\prime}}-2\Omega)^{2}]}\biggr{]}\ ,$ (S25)
and
$\mathcal{I}_{JK}^{(m^{\prime})}$ $=\int_{-\infty}^{\infty}dt_{2}\int_{-\infty}^{t_{2}}dt_{1}\,e^{-\frac{t_{2}^{2}}{2t_{d}^{2}}}\cos(\Omega t_{2})\,e^{-\frac{t_{1}^{2}}{t_{d}^{2}}}[1+\cos(2\Omega t_{1})]\,e^{i\omega_{m^{\prime}0}t_{1}}e^{i\omega_{mm^{\prime}}t_{2}}$
$\approx\frac{\pi t_{d}}{2}\,e^{-\frac{t_{d}^{2}}{4}\omega_{m^{\prime}0}^{2}}\left[e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}-\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}+\Omega)^{2}}\right]$
$\quad+\frac{\pi t_{d}}{4}\left[e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}-2\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{4}(\omega_{m^{\prime}0}+2\Omega)^{2}}\right]\left[e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}-\Omega)^{2}}+e^{-\frac{t_{d}^{2}}{2}(\omega_{mm^{\prime}}+\Omega)^{2}}\right]\ ,$ (S26)
denoting resonant processes at different frequencies. For instance, the $JJ$ case, leads to one-photon resonances at $\omega_{m^{\prime}0}\approx\pm\Omega$ and $\omega_{mm^{\prime}}\approx\pm\Omega$. The Gaussian envelope sharply restricts these resonances, ensuring that only states close to these conditions contribute significantly. The $KK$ case allows for two-photon processes or virtual zero-energy components, with resonances occurring at $\omega_{m^{\prime}0},\omega_{mm^{\prime}}\approx 0,\pm 2\Omega$, and the mixed case $JK$ couples one-photon and two-photon contributions, describing interference between different harmonics.
2 Relevant Matrix elements
The perturbative analysis above emphasizes the role of the current and kinetic energy operators in the fate of dynamically accessing target states. In the main text, we have focused on the $\hat{J}_{x}$ matrix elements in the eigenbasis of the unperturbed Hamiltonian. Now, we extend this analysis by showing the corresponding $y$-direction in Fig. S1(b). Unlike the $x$-direction case, the matrix $\langle m|\hat{J}_{y}|n\rangle$ is much sparser and does not exhibit significant matrix elements of the form $\langle 2|\hat{J}_{y}|0\rangle$, thus precluding first-order excitation of the target state $|2\rangle$, nor simultaneously large $\langle 2|\hat{J}_{y}|m^{\prime}\rangle$ and $\langle m^{\prime}|\hat{J}_{y}|0\rangle$, preventing relevant second-order current-current excitation terms. This justifies the choice of $x$-polarized pulses we implement in the main text.
Figure S1: The absolute value of the current matrix elements in the eigenstate basis of the Hamiltonian, $|\langle m|\hat{J}_{\mu}|n\rangle|$, focusing on the 30 states with lowest energy, for $\mu=x$ in (a) and $\mu=y$ in (b). The dashed lines highlight the target state $|2\rangle$ used in the main text.
Additionally, a strong indication that a second-order process is responsible for the excitation of eigenstate $|2\rangle$ is given by the analysis of the matrix elements of $\hat{J}_{x}^{2}$ in the eigenbasis of $\hat{\cal H}$. As Eqs. (S12) and (S24) indicate, relevant terms as $\propto\sum_{m^{\prime}}\langle m|\hat{J}_{x}|m^{\prime}\rangle\langle m^{\prime}|\hat{J}_{x}|0\rangle=\langle m|\hat{J}_{x}^{2}|0\rangle$ need to be large for this type of contribution to be important. Indeed, as Fig. S2(b) shows, the most significant matrix elements of $\hat{J}_{x}^{2}$ in the low-lying spectrum are in between states $|2\rangle$ and $|0\rangle$.
Finally, contributions for a first-order transition between $|0\rangle$ and $|2\rangle$ in the time-dependent perturbation can be accomplished by using the kinetic energy operator, see Eq. (S22), with matrix elements proportional to $\langle m|\hat{K}_{x}|0\rangle$. As Fig. S2(a) indicates, $|\langle 2|\hat{K}_{x}|0\rangle|$ is the largest in the low-lying spectrum. Nonetheless, because the coefficient in the first-order transition amplitude is $A^{2}(t)$ for the kinetic energy operator, unlike $A(t)$ for the current one, this dictates that such kinetic-driven excitation is not significantly favored for small pulse amplitudes.
Figure S2: The absolute value of the matrix elements of the $x$-direction kinetic energy (a) and $x$-direction squared-current (b) operators in the eigenstate basis of the Hamiltonian, focusing on the 12 states with lowest energy. Here, the diagonal elements are set to zero for visualization (they are typically much larger). The dashed lines highlight the target state $|2\rangle$ used in the main text.
3 Density dynamics
The main text focuses on the melting of the charge order reversal with the pulses, releasing a suppressed zero-momentum occupancy. Meanwhile, we can also investigate the fate of the imposed stripes by the local chemical potential $V_{0}$ under photoirradiation. Figure S3 shows the dynamics of the densities for the four inequivalent sites (under periodic boundary conditions) through photoirradiation with the same conditions as Figs. 1 and 3 of the main text, i.e., under maximization of the late-time overlap $|\langle 2|\psi(t_{\rm max})\rangle|$.
Figure S3: Dynamics for the density resolved in the inequivalent sites under periodic boundary conditions – inset depicts the lattice with inequivalent sites annotated from 0 to 3. Photoirradiation conditions are the same of Fig. 1 in the main text: $\Omega=1.713\ t_{h}$, $A_{0}=0.62$ and $t_{d}=6.5t_{h}^{-1}$, for Hamiltonian parameters $V/t_{h}=4$ and $V_{0}/t_{h}=5$.
Before the pump, the interstripe density is roughly the same (in particular $\langle\hat{n}_{1}\rangle=\langle\hat{n}_{3}\rangle$ due to symmetry considerations), and much larger than the one for stripe sites ($\langle\hat{n}_{0}\rangle\simeq 0.045$). After photoirradiation, although $\pi$-phase shift is now absent, the density imbalance between stripe and interstripe sites is substantially reduced, but still significant, indicating that the externally imposed stripes still survive the dynamical perturbation, albeit with a smaller strength.
Figure S4: Dependence of the long-time average of the zero momentum occupancy in the pulse parameters $A_{0}$, pulse amplitude, and $t_{d}$, pulse width. Here the pulse frequency is the one used in Figs. 1 and 3 of the main text, $\Omega=1.713\ t_{h}$. The white-edge marker describes the pair of parameter values ($A_{0},t_{d}$) used in the main text.
4 Maximizing zero-momentum occupancy
In the main text, we establish what pump parameters, namely $A_{0}$, $t_{d}$, and $\Omega$, maximize the overlap with the target state, the equilibrium eigenstate $|2\rangle$ of $\hat{\cal H}$, as emphasized in Fig. 4(a). We now show that this set of parameters also maximizes the late-time averaged zero momentum occupancy $\langle\hat{n}_{{\bf k}=0}\rangle$ sufficiently after the photoirradiation. Figure S4 shows the time average of the zero momentum occupancy at long times ($t\in[4.5t_{d}-4t_{h}^{-1},4.5t_{d}]$) as a function of the pump amplitude and width for the frequency $\Omega=1.713\ t_{h}$ — it is clear that condensation is enhanced in the same range of parameters in which the overlap $|\langle\psi(t\gg t_{d})|2\rangle|^{2}$ is maximized.
Figure S5: Dependence on $A_{0}$ and $t_{d}$ of the long-time weights of equilibrium states in $|\psi(t)\rangle$, for $|m=0\rangle$ (a), $|2\rangle$ (b), and $|4\rangle$ (c) — the marker assigns the parameters $(A_{0},t_{d})=(0.34,10.8)$ resulting in the maximum overlap with the target state $|2\rangle$. (d) Pulse representation for these parameter settings. Dynamics of the connected density correlations across a stripe (e), and the zero momentum occupancy (f). (g) Selected overlaps, by higher values in the dynamics, of the time-evolving state during the photoirradiation process. Data in the plots employ the second-order resonant photoirradiation $\Omega=1.491\ t_{h}=\omega_{42}$.
5 Other intermediate states
While analyzing the matrix elements of the current operator $\hat{J}_{x}$ [Fig. S1], we notice that other states can, in principle, serve as intermediate states for a dynamical transition $|0\rangle\to|2\rangle$. For instance, considering $|m^{\prime}=4\rangle$, we notice that $|\langle 2|\hat{J}_{x}|4\rangle|=0.24$ and $|\langle 4|\hat{J}_{x}|0\rangle|=1.73$ — the largest among the lowest 30 eigenpairs of $\hat{\cal H}$. Thus, using a similar prescription of the main text for the case $|m^{\prime}=6\rangle$, resonant and energy conservation conditions lead, respectively, to $E_{4}-E_{0}\approx n_{1}\Omega$ and $E_{4}-E_{0}=(n_{1}+n_{2})\Omega$. Subtracting these, we obtain $\omega_{42}\simeq-n_{2}\Omega$. As in the main text, if we consider the $|4\rangle\to|2\rangle$ dynamical process as a one-photon emission ($n_{2}=-1$), we can set the pump frequency $\Omega=\omega_{42}\simeq 1.491$. Now combining with the resonant condition, yields $\omega_{40}/\Omega\approx n_{1}=4.475/1.491\simeq 3$, indicating a similar three quantum absorption scenario.
Besides that, additional tuning of the pulse amplitudes and time-width is needed in order to target the desired state, $|2\rangle$, properly. This is accomplished by analyzing the long-time weight of selected equilibrium eigenstates $|m\rangle$ in $|\psi(t)\rangle$, as shown in Fig. S5(a)–(c). Maximimization of $|\langle 2|\psi(t_{\rm max})\rangle|^{2}$ is achieved, with $\Omega=\omega_{42}=1.491\ t_{h}$ at $A_{0}=0.34$ and $t_{d}=10.8\ t_{h}^{-1}$. Focusing on these parameters [see Fig. S5(d) for the pulse representation], we show the dynamics of both the connected density correlations [Fig. S5(e)] as well as the zero momentum occupancy [Fig. S5(f)]. In this case, the enhancement of the average $n_{{\bf k}=0}$ of about $26\%$ after the pulse is masked by the substantial oscillations that well encompass values below that of the equilibrium regime. Moreover, despite the reduction of the $\pi$-phase density correlations across a stripe, their melting is not achieved in these settings. This suggests that the dynamical addressing of the target state was not optimized.
Indeed, as the dynamics of the overlaps in Fig. S5(g) displays, the weight of $|2\rangle$ in $|\psi(t)\rangle$ is yet small in comparison with the one with the initial state, the ground-state of $|0\rangle$. As a consequence, the targeting is only partial, in contrast to that obtained in the main text.